OpenStudy (anonymous):

The ratio of men to women working for a company is 7 to 5 . If there are 288 employees total, how many men work for the company?

OpenStudy (anonymous):

@sangya21 @quickstudent @uri @mathmath333

OpenStudy (anonymous):

@hhelpplzzzz @rianiscool @eliassaab @kaylamarie081 help

OpenStudy (anonymous):

let men be x 7/5 = x/(288-x) 7(288-x) =5x 2016 -7x = 5x 2016 = 12x 2016 / 12= x x = 168

OpenStudy (anonymous):

i tried that , i thought it was wrong but it was rght :) Brian is going to rent a truck for one day. There are two companies he can choose from, and they have the following prices. Company A charges $91 and allows unlimited mileage. Company B has an initial fee of $55 and charges an additional $0.90 for every mile driven. For what mileages will Company A charge less than Company B? Use m for the number of miles driven, and solve your inequality for m

OpenStudy (anonymous):

@sangya21 @mathmath333

OpenStudy (anonymous):

91< 55+0.90m i think so but not sure

OpenStudy (mathmath333):

\(\Large 91< 55+0.9n\)